OpenStudy (anonymous):

Find the LCM of 21y^2) + 42y and 3y^2 + 9y + 6

OpenStudy (ybarrap):

21y^2 + 42y = 21y(y + 2) 3y^2 + 9y + 6 = 3*(y + 1)(y + 2) LCM = 21y(y + 2)*3*(y+1)(y+2)/[3*(y + 2)] = 21y*(y+1)(y+2) Ans. 21y*(y+1)(y+2)